OpenStudy (anonymous):

solve by substitution method 3x+7y=5 6x+4y=10

OpenStudy (anonymous):

(5/2-3x/2)=y 3x+7((5/2-3x/2))=5 x=5/3 y=5/2-3/2(5/3) = 5/2-5/2=0 so x=5/3 and y = 0

OpenStudy (anonymous):

thank you!!